OpenStudy (anonymous):

what are the solutions to g(x)=lnx-5 walk through please?

OpenStudy (zzr0ck3r):

you have just listed the function. Do you mean what are the solutions to \(f(x) =0\)?

OpenStudy (zzr0ck3r):

I am going to assume this is what you mean. note \(\ln(x)-5=0\implies \ln(x) = 5 \) Also note that \[\log_a(b)=c \iff a^c = b\] Can you figure it out now?

OpenStudy (anonymous):

e=5...?

OpenStudy (zzr0ck3r):

\(\log_a(b)=c \iff a^c = b \) \(\log_e(x)=5 \iff ?= x\)

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

e^5=x

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

so...x=148.413 or can I leave it as e^5=x?

OpenStudy (zzr0ck3r):

I would leave it. The first way is an approximation. The second way is the exact answer.

OpenStudy (anonymous):

haha, ok thanks(:

OpenStudy (zzr0ck3r):

Unless you are told otherwise....

OpenStudy (zzr0ck3r):

np

OpenStudy (zzr0ck3r):

I think of it like this \(\log_a(x)=y \iff x = a^y \)